OpenStudy (anonymous):

how do you get rid of the exponent in x^4=1?

OpenStudy (anonymous):

on order to solve for x

OpenStudy (ivettef365):

by sqrt on both sides

OpenStudy (anonymous):

even though its x^4 and not x^2?

OpenStudy (anonymous):

x^4 = 1 x^2 = -1 or 1 x = -1,1,-i,i

OpenStudy (anonymous):

so if i square root both sides i get x= 1?

OpenStudy (anonymous):

No if you sqrt both sides you get $x^2={\plusminus}1$

OpenStudy (anonymous):

how do i get the final answer?

OpenStudy (anonymous):

Then square root both sides again. The sqrt of 1 is \[{\plusminus}1\] The sqrt of -1 is \[{\plusminus}i\] But of you guys haven't learner about \[i\] yet, then just say \[{plusminus}1\]. But if you do know about I, it would be useful to learn about DeMoivre's Theorem.

OpenStudy (anonymous):

okay, thank you so much!

OpenStudy (zarkon):

the plus minus \(\pm\) is just "\pm" in\(\LaTeX\)

OpenStudy (anonymous):

Haha, thanks, I was wondering why that came out so funky. Also, is there any way to put it unlike?

OpenStudy (anonymous):

*inline

OpenStudy (zarkon):

\ ( latex here \ ) but don't put a space between \ and (

OpenStudy (anonymous):

Ok, \(thanks\)

OpenStudy (anonymous):

what okay, I have a question. So I was trying to find x in order to solve this problem: Find the solutions to 0 ≥ x4 – 1 by graphing y ≥ x4 – 1 by hand. So now I'm plugging in \[\pm i\] into y ≥ x4 – 1 but I dont know how to lol sorry

OpenStudy (phi):

I think they expect only the real solutions to be plotted

OpenStudy (anonymous):

Thats what I thought, so how do I solve it without using i

OpenStudy (phi):

The want you to graph y= x^4 -1 can you do that ?

OpenStudy (anonymous):

Thats it, I think

OpenStudy (phi):

It does not look like you did it by hand, but that is the curve now identify where y ≤ 0 (where y = x^4 – 1 )

OpenStudy (anonymous):

i did it by hand on my paper I just didnt know how else to show it to you lol thats it with y </= 0

OpenStudy (phi):

the curve is ≤ 0 between -1 ≤ x ≤ 1

OpenStudy (phi):

Find the solutions to 0 ≥ x^4 – 1 is -1 ≤ x ≤ 1

OpenStudy (anonymous):

alright, thank you!!